OpenStudy (anonymous):

Given\[A=\left[\begin{matrix}2 & 1 & 3 \\ 0 & -1 & 4 \\ 0&0&0\end{matrix}\right]\]Determine the eigenvalues and who that the corresponding eigenspaces are dimension 1 and are generated by the eigenvectors\[\left(\begin{matrix}1 \\ 0\\0\end{matrix}\right),\left(\begin{matrix}1 \\ -3\\0\end{matrix}\right),\left(\begin{matrix}-7 \\ 8\\2\end{matrix}\right)\]

OpenStudy (dumbcow):

Find determinate of (A-lambda) and set equal to 0 \[\det\left[\begin{matrix}2-L& 1&3 \\ 0 & -1-L&4\\0&0&-L\end{matrix}\right] = L(2-L)(1-L)=0\] \[L = 0,1,2\]

OpenStudy (dumbcow):

for eigenvalue, L=0 \[\left[\begin{matrix}2& 1&3 \\ 0 & -1&4\\0&0&0\end{matrix}\right]\left[\begin{matrix} x \\ y\\z\end{matrix}\right]= \left[\begin{matrix} 0 \\ 0\\0\end{matrix}\right]\]

OpenStudy (dumbcow):

solving system yields... \[x = -(7/2)z\] \[y = 4z\] If z = 2, then x = -7, y = 8 therefore eigenvector (-7,8,2) goes with eigenvalue L=0 do the other 2 eigenvalues the same way

OpenStudy (dumbcow):

here is a reference http://en.wikipedia.org/wiki/Eigenvalues_and_eigenvectors